OpenStudy (writerforever18):

Consider the line y=3/5x-1 Find an equation of the line that is parallel to this and passes through the point (3, 3) Find an equation of the line that is perpendicular to this and passes through the point (3, 3). Please help!

OpenStudy (anonymous):

In order to find the equation of the line that is perpendicular to y = 3/5 x we have to know what makes a line perpendicular to another. Do you know? It has a inverse slope, so for this problem, the perpendicular line would have slope -5/3. To find the equation of the line with -5/3 slope and passes through point (5,3) we must use the point-slope form of an equation. y - y1 = m(x- x1) Therefore if we substitute -5/3 for the m and 3 for y1 and 5 for the x1 we get y-3 = -5/3(x-5). Then simplify to slope intercept form. y-3 = -5/3x +25/3 y-3+3 = -5/3x +25/3 +3 y = -5/3x +(25+9)/3 y = -5/3x + 34/3 y = -5/3x + 11 1/3 Hope this helps, Pam W.

OpenStudy (writerforever18):

So the perpendicular slope is y=5/3x+11 and 1/3?

OpenStudy (writerforever18):

Do you know what 11 1/3 is as a fraction?